OpenStudy (anonymous):

Physics help please!! Will give medal! A rock dropped from a high platform is moving at 24 m/s downward when it strikes the ground. Ignore air resistance. How fast was the rock moving when it had fallen only one-fourth of the distance to the ground? 18.0 m/s 12.0 m/s 8.0 m/s 6.0 m/s

OpenStudy (anonymous):

@lonnie455rich

OpenStudy (anonymous):

I'm thinking it's either 6 or 8

OpenStudy (anonymous):

final velocity is 24 m/s and gravity is 9.8 m/s so what equations do you have for these two

OpenStudy (anonymous):

v^2 = u^2 + 2ax is what my teacher told us to use I think

OpenStudy (anonymous):

define your variables.

OpenStudy (anonymous):

right sorry! x=displacement u= initial velocity v = final velocity t= time and a=acceleration which I suppose is 9.8m/s?

OpenStudy (anonymous):

you know your initial velocity is 0 correct.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now what do I do? @lonnie455rich

OpenStudy (anonymous):

plug and chug to dind displacement

OpenStudy (anonymous):

then plug that displacement in to find final velocity at 1/4 the displacement

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

24 = 0^2 + 2(9.8)(d)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ya. that is it. unless you are supposed to square the final velocity. i cant remember

OpenStudy (anonymous):

i think you square it

OpenStudy (anonymous):

yeah sorry I meant 24^2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

sorry i have to go but @petiteme can help. you should have it figured out, but just in case.

OpenStudy (anonymous):

@petiteme

OpenStudy (anonymous):

is it 29.4 approximately?

OpenStudy (anonymous):

I mean 7.4?

OpenStudy (anonymous):

@petiteme

OpenStudy (petiteme):

Did you solve it?

OpenStudy (anonymous):

yes 7.4 is what I got

OpenStudy (anonymous):

@petiteme

OpenStudy (anonymous):

hello????